OpenStudy (anonymous):

How do you factor x^4+6x^3+8x+48?

OpenStudy (anonymous):

use grouping: factor our x^3 from the first two terms, and 8 from the second, like this: x^3(x+6)+8(x+6) so then this equals: (x^3+8)(x+6). Then the first binomial is a sum of cubes (2^3=8) so we factor that using the sum of cubes rule: (a^3+b^3)=(a+b)(a^2+ab+b^2) so the final answer is: (x+6)(x+2)(x^2+2x+4)

OpenStudy (anonymous):

Thank you :)